# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ Abstract
Probabilistic electricity price forecasting (PEPF) is a key task for market participants in short-term electricity markets. The increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. We present an online, multivariate, regularized distributional regression model, allowing for the modeling of all distribution parameters conditional on explanatory variables. Our approach is based on the combination of the multivariate distributional regression and an efficient online learning algorithm based on online coordinate descent for LASSO-type regularization. Additionally, we propose to regularize the estimation along a path of increasingly complex dependence structures of the multivariate distribution, allowing for parsimonious estimation and early stopping. We validate our approach through one of the first forecasting studies focusing on multivariate probabilistic forecasting in the German day-ahead electricity market while using only online estimation methods. We compare our approach to online LASSO-ARX-models with adaptive marginal distribution and to online univariate distributional models combined with an adaptive Copula. We show that the multivariate distributional regression, which allows modeling all distribution parameters – including the mean and the dependence structure – conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only and keeping a static/unconditional dependence structure. Additionally, online estimation yields a speed-up by a factor of 80 to over 400 times compared to batch fitting.
Keywords: online learning, GAMLSS, LASSO, covariance estimation, cholesky-decomposition, low-rank approximation, multivariate distributional regression, probabilistic electricity price forecasting, day-ahead electricity market, EPF
@@ 1 Introduction
Short-term electricity markets play a key role in the integration of renewable energy sources and flexible generation in the electricity system. In Germany, the day-ahead auction is the major venue for physically delivered electricity. Trading volumes have grown with the increase of renewable generation capacity. To optimize decision-making and bidding strategies, market participants need accurate price forecasts. Additionally, electricity prices are multivariate time series characterized by high volatility, positive and negative spikes and skewness. Therefore, research and industry have moved towards probabilistic electricity price forecasting (PEPF) to account for their stochastic nature (see e.g. Nowotarski and Weron, 2018; Dexter Energy, 2024). However, the multivariate dimension of electricity price time series has received little attention for PEPF so far, while being of high importance for market participants in the context of the optimization of flexible assets and portfolio management (Löhndorf and Wozabal, 2023; Peña et al., 2024; Beykirch et al., 2022, 2024). At the same time, the increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. This work presents an online, multivariate distributional regression model, which we apply for probabilistic day-ahead electricity price forecasting in Germany. Our work is among the first to treat the 24-dimensional hourly electricity prices as multivariate distribution and the first to treat the problem in a strict online estimation setting, which makes the complex, high-dimensional distributional learning problem feasible on standard laptops. Our results show that modeling the dependence structure improves forecasting performance significantly compared to univariate approaches.
@@
The need for multivariate PEPF
The literature on PEPF has evaluated a wide range of different statistical and machine learning methods, such as quantile regression, ARX-GARCH models (Nowotarski and Weron, 2018; Billé et al., 2023; Marcjasz et al., 2023), conformal prediction methods, (see e.g. Kath and Ziel, 2021; Zaffran et al., 2022; Brusaferri et al., 2024a; Lipiecki et al., 2024), distributional regression and neural network approaches (e.g. Brusaferri et al., 2024b; Marcjasz et al., 2023; Hirsch et al., 2024; Ziel et al., 2021). However, these works treat each delivery hour as independent, univariate time series as in Ziel and Weron (2018). Let us motivate the need for multivariate probabilistic forecasting approaches for the day-ahead electricity price by two simple plots. Figure 2 shows a time series plot for the 24 hourly day-ahead electricity prices in Germany. The left panel shows each delivery hour as individual, daily series, emphasizing the daily co-movement. The right panel shows the cross-section, i.e. the daily shape for the first 180 days of 2017. The temporal correlations along the dimension of the delivery hours $h=0,...,23$ is clearly visible. Additionally, Figure 1 shows the correlation matrix of the raw electricity prices, but also the residual correlation for standard LASSO-ARX models for the electricity price. We see a strong, statistically significant remaining residual cross-correlation, indicating that the resulting marginal error distributions, which are conditional on the mean, are not independent. On top of the statistical motivation, Beykirch et al. (2022, 2024) clearly describe the need for predicting joint distributions for the optimization of schedules and bidding curves in energy markets, further examples are provided by Peña et al. (2024); Löhndorf and Wozabal (2023).
Figure 2: Time series plot for day-ahead electricity prices $P_{d,h}$ in Germany. In the left panel, each color corresponds to one delivery hour $h=0,...,23$. The blue dotted line marks the split between test and training data set. The gray area corresponds to the time of the right panel, which shows the same data along the dimension of the delivery hour, where each line represents a delivery day $d$. The high volatility, occasional positive and negative spikes and co-movement of electricity prices are visible.
Literature on multivariate PEPF
@@
Work on multivariate probabilistic forecasting for day-ahead electricity prices are sparse in the literature and the majority of the existing works, e.g. Maciejowska and Nitka (2024); Berrisch and Ziel (2024); Han (2023); Mashlakov et al. (2021) and Agakishiev et al. (2025), does not evaluate multivariate scoring rules such as the VS, DSS or ES, but focuses on the evaluation of the marginals of the multivariate distribution through the CRPS. This reduces the problem to modeling 24 marginal distributions, taking only lagged cross-information into account. To the best knowledge of the author, only two studies truly model and evaluate the multivariate dependence structure. First, Janke and Steinke (2020) approach the issue through implicit generative Copula models. Grothe et al. (2023) employ the Schaake shuffle, a post-processing method for point forecasts. On the contrary, in the fields of probabilistic weather, renewable production (Bjerregård et al., 2021; Sørensen et al., 2022; Kolkmann et al., 2024) and probabilistic load forecasting (Gioia et al., 2022; Browell et al., 2022) truly multivariate forecasting approaches have gained more attention.
Distributional Regression
@@
The goal of distributional regression or “regression beyond the mean” (Kneib et al., 2023; Klein, 2024) is modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditional on explanatory variables. The most prominent model in this regard is the original GAMLSS (Generalized Additive Model for Location, Scale and Shape Rigby and Stasinopoulos, 2005), of which numerous extensions have been developed over the last years (Kock and Klein, 2023; Kneib et al., 2023; Muschinski et al., 2022) and distributional deep neural networks (DDNN, e.g. Klein et al., 2021, 2023; Rügamer et al., 2024). Through the direct modeling of the variable’s distribution, this method is well suited for the generation of probabilistic forecasts and has been successfully applied in energy markets (Muniain and Ziel, 2020; Gioia et al., 2022; Serinaldi, 2011; Brusaferri et al., 2024b; Marcjasz et al., 2023).
Online Learning
@@
For environments with large amounts of continuously incoming data, such as energy markets, online learning describes the task of updating the model given new data, without falling back on previous samples. Formally, in the strict online setting, after having seen $N$ samples of our data set, we fit a model, predict for step $N+1$. Subsequently, we receive the realized values for $N+1$ and update our model, taking into account only the new row $N+1$. This approach allows an efficient processing of high-velocity data and results in greatly decreased computational effort. The principle is outlined as well in Figure 6. Online learning for LASSO-regularized regression for the mean has been introduced in Angelosante et al. (2009, 2010) and Messner and Pinson (2019). Univariate approaches suitable for probabilistic forecasting based stochastic gradient descent have been developed for specific distributions, (see e.g. Pierrot and Pinson, 2021), conformal prediction (see e.g. Zaffran et al., 2022; Brusaferri et al., 2024a) and the generic online distributional regression in Hirsch et al. (2024), however, in the multivariate case, the literature remains sparse and focused on unconditional distributions and Copulae (see e.g. Dasgupta and Hsu, 2007; Zhao et al., 2022; Landgrebe et al., 2020).
Contributions
@@
We add to the literature by presenting a generic, online, regularized, multivariate distributional regression model, allowing to model all distribution parameters conditional on explanatory variables and validate the approach in a forecasting study for the day-ahead electricity market in Germany. Our paper is the first to tackle the issue of truly multivariate probabilistic energy in a strict online estimation setting. In detail, our contributions include
  • We present the online, multivariate distributional regression model based on the combination of the multivariate distributional regression (Muschinski et al., 2022; Kock and Klein, 2023; Gioia et al., 2022) and implement an efficient online learning algorithm based on the univariate work by Hirsch et al. (2024).
  • We propose a regularized estimation, using both, LASSO for each individual distribution parameter, but also a path-based estimation along increasingly complex dependence structures in the multivariate distribution, allowing for parsimonious estimation and early stopping.
@@
  • Our case study explores the multivariate normal and multivariate $t$-distribution for the joint distribution of spot electricity prices, compared to both online LASSO-ARX-models (see e.g. Nowotarski and Weron, 2018) with constant marginal distributions, but also to the online univariate distributional model by Hirsch et al. (2024) combined with a Copula approach. We show that the multivariate distributional regression, which allows modeling all distributional parameters, i.e. the mean, but also the dependence structure, conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only respectively keeping a static/unconditional dependence structure.
@@
  • We providing a high-performing Python implementation using just-in-time compilation and providing a familiar, scikit-learn-like API to facilitate the usage of our package for other researchers. Reproduction code can be found in the GitHub repository at https://github.com/simon-hirsch/online-mv-distreg. We plan to integrate our code in the ROLCH package by Hirsch et al. (2024).
Structure of the paper
@@
The remainder of the paper is structured as usual: The following main Section 2 introduces the multivariate, online, regularized distributional regression model. Sections 3 introduce the forecasting study and the used data and 4 presents our results. Section 5 concludes the paper.
@@ 2 Online Multivariate Distribution Regression
We start the exposition on distributional regression by building from the univariate, batch case onto the multivariate setting and subsequently moving to the online setting. The following section introduces the necessary notation, briefly reviews online learning approaches for LASSO-regularized estimation present and the framework of distributional regression in a rather general way. Subsequently, we discuss possible options for achieving (almost) unconstrained parametrization for the scale and precision matrix of multivariate distributions (Section 2.2). Iteratively reweighed least squares for the estimation is introduced in Section 2.3 and the online algorithm is introduced in Section 2.4. Lastly, we discuss path-based regularization and early stopping in Section 2.5.
2.1 Preliminaries and Setting
Notation

### v2

@@ Abstract
Probabilistic electricity price forecasting (PEPF) is vital for short-term electricity markets, yet the multivariate nature of day-ahead prices — spanning 24 consecutive hours — remains underexplored. At the same time, real-time decision-making requires methods that are both accurate and fast. We introduce an online algorithm for multivariate distributional regression models, allowing an efficient modelling of the conditional means, variances, and dependence structures of electricity prices. The approach combines multivariate distributional regression with online coordinate descent and LASSO-type regularization, enabling scalable estimation in high-dimensional covariate spaces. Additionally, we propose a regularized estimation path over increasingly complex dependence structures, allowing for early stopping and avoiding overfitting. In a case study of the German day-ahead market, our method outperforms a wide range of benchmarks, showing that modeling dependence improves both calibration and predictive accuracy. Furthermore, we analyse the trade-off between predictive accuracy and computational costs for batch and online estimation and provide an high-performing open-source Python implementation in the ondil package.
keywords:
online learning , multivariate distributional regression , probabilistic electricity price forecasting , LASSO regularization , day-ahead electricity market
[…]
@@ 1 Introduction
Short-term electricity markets play a key role in the integration of renewable energy sources and flexible generation in the electricity system. In Germany, the day-ahead auction is the major venue for physically delivered electricity. Trading volumes have grown with the increase of renewable generation capacity. To optimize decision-making and bidding strategies, market participants need accurate price forecasts. Since electricity prices are characterized by high volatility, positive and negative spikes and skewness, research and industry have moved towards probabilistic electricity price forecasting (PEPF) to account for their stochastic nature (see e.g. Nowotarski and Weron, 2018; Dexter Energy, 2024). With 24 prices per day, electricity prices are multivariate time series with a potentially complex dependency structure. However, the multivariate dimension of electricity price time series has received little attention for PEPF so far, while being of high importance for market participants in the context of the optimization of flexible assets and portfolio management (Löhndorf and Wozabal, 2023; Peña et al., 2024; Beykirch et al., 2022, 2024). At the same time, the increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. This work presents an online, multivariate distributional regression model, which we apply for probabilistic day-ahead electricity price forecasting in Germany. Our work is among the first to treat the 24-dimensional hourly electricity prices as multivariate distribution and the first to treat the problem in a strict online estimation setting, which makes the complex, high-dimensional distributional learning problem feasible on standard laptops. Our results show that modeling the dependence structure improves forecasting performance significantly compared to univariate approaches.
Figure 1: Time series plot for day-ahead electricity prices $P_{d,h}$ in Germany. In the left panel, each color corresponds to one delivery hour h=0,,23h=0,...,23. The blue dotted line marks the split between test and training data set. The gray area corresponds to the time of the right panel, which shows the same data along the dimension of the delivery hour, where each line represents a delivery day $d$. The high volatility, occasional positive and negative spikes and co-movement of electricity prices are visible.
Mean $\mu$ Std $\sigma$ MAD Min Max
Pre-2021 34.6 16.6 8.8 -130.1 200.0
2021 96.8 73.7 26.3 -69.0 620.0
2022 235.5 142.8 86.2 -19.0 871.0
2023 95.2 47.6 23.1 -500.0 524.3
Table 1: Summary statistics for day-ahead electricity prices $P_{d,h}$ in Germany for the years 2015 to 2023. The mean $\mu$, standard deviation $\sigma$, median absolute deviation (MAD), minimum and maximum are given.
The literature on PEPF has evaluated a wide range of different statistical and machine learning methods, such as quantile regression, ARX-GARCH models (Nowotarski and Weron, 2018; Billé et al., 2023; Marcjasz et al., 2023), conformal prediction methods, (see e.g. Kath and Ziel, 2021; Zaffran et al., 2022; Brusaferri et al., 2024a; Lipiecki et al., 2024), distributional regression and neural network approaches (e.g. Brusaferri et al., 2024b; Marcjasz et al., 2023; Hirsch et al., 2024; Ziel et al., 2021). However, these works treat each delivery hour as independent, univariate time series as in Ziel and Weron (2018). Let us motivate the need for multivariate probabilistic forecasting approaches for the day-ahead electricity price by two simple plots. Figure 1 shows a time series plot for the 24 hourly day-ahead electricity prices in Germany. The left panel shows each delivery hour as individual, daily series, emphasizing the daily co-movement. The right panel shows the cross-section, i.e. the daily shape for the first 180 days of 2017. The temporal correlations along the dimension of the delivery hours h=0,,23h=0,...,23 is clearly visible. We also like to highlight the high volatility during the winter 2021 and the Russian invasion of Ukraine 2022 and 2023. The long-run average prices have shifted from roughly 35 EUR/MWh in prior to 2021 to over 235 EUR/MWh in 2022. In the same reign, the standard deviation has increased from roughly 16 EUR/MWh to over 85 EUR/MWh. Summary statistics are given in Table 1, where the years 2021, 2022 and 2023 are shown separately to highlight the changes in the market. Additionally, Figure 2 shows the correlation matrix of the raw electricity prices (lower triangular) and the residual correlation for standard LASSO-ARX models for the electricity price (upper triangular). We see a strong, statistically significant remaining residual cross-correlation, indicating that the resulting marginal error distributions, which are conditional on the mean, are not independent. On top of the statistical motivation, Beykirch et al. (2022, 2024) clearly describe the need for predicting joint distributions for the optimization of schedules and bidding curves in energy markets, further examples are provided by Peña et al. (2024); Löhndorf and Wozabal (2023).
@@
Work on multivariate probabilistic forecasting for day-ahead electricity prices are sparse in the literature and the majority of the existing works, e.g. Maciejowska and Nitka (2024); Berrisch and Ziel (2024); Han (2023); Mashlakov et al. (2021) and Agakishiev et al. (2025), do not evaluate multivariate scoring rules such as the VS, DSS or ES, but focus on the evaluation of the marginals of the multivariate distribution through the CRPS. This reduces the problem to modeling 24 marginal distributions, taking only lagged cross-information into account. To the best knowledge of the author, only two studies truly model and evaluate the multivariate dependence structure. First, Janke and Steinke (2020) approach the issue through implicit generative Copula models. Grothe et al. (2023) employ the Schaake shuffle, a post-processing method for point forecasts. On the contrary, in the fields of probabilistic weather, renewable production (Bjerregård et al., 2021; Sørensen et al., 2022; Kolkmann et al., 2024) and probabilistic load forecasting (Gioia et al., 2022; Browell et al., 2022) truly multivariate forecasting approaches have gained more attention.
@@
The goal of distributional regression or “regression beyond the mean” (Kneib et al., 2023; Klein, 2024) is modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditional on explanatory variables. The most prominent model in this regard is the original Generalized Additive Model for Location, Scale and Shape (GAMLSS, Rigby and Stasinopoulos, 2005), of which numerous extensions have been developed over the last years (Kock and Klein, 2023; Kneib et al., 2023; Muschinski et al., 2022) and distributional deep neural networks (DDNN, e.g. Klein et al., 2021, 2023; Rügamer et al., 2024). Through the direct modeling of the variable’s distribution, this method is well suited for the generation of probabilistic forecasts and has been successfully applied in energy markets (Muniain and Ziel, 2020; Gioia et al., 2022; Serinaldi, 2011; Brusaferri et al., 2024b; Marcjasz et al., 2023). A drawback of fully distributional models is the computational effort and the need for efficient estimation approaches for (multivariate) distributional regression models has been recognized by Umlauf et al. (2025) and Gioia et al. (2025), who propose efficient batch estimation approaches.
@@
For environments with large amounts of continuously incoming data, such as energy markets, online learning describes the task of updating the model given new data, without falling back on previous samples. Formally, in the strict online setting, after having seen $N$ samples of our data set, we fit a model, predict for step $N+1$. Subsequently, we receive the realized values for $N+1$ and update our model, taking into account only the new row $N+1$. This approach allows an efficient processing of high-velocity data and results in greatly decreased computational effort. The principle is outlined as well in Figure 6. Online learning for LASSO-regularized regression for the mean has been introduced in Angelosante et al. (2009, 2010) and Messner and Pinson (2019). Univariate approaches suitable for probabilistic forecasting based stochastic gradient descent have been developed for specific distributions, (see e.g. Pierrot and Pinson, 2021), conformal prediction (see e.g. Zaffran et al., 2022; Gibbs and Candes, 2021; Gibbs and Candès, 2024) and the generic online distributional regression in Hirsch et al. (2024). However, in the multivariate case, the literature remains sparse and focused on unconditional distributions and Copulae (see e.g. Dasgupta and Hsu, 2007; Zhao et al., 2022; Landgrebe et al., 2020).
@@
We add to the literature by presenting a generic, online, regularized, multivariate distributional regression model, allowing to model all distribution parameters conditional on explanatory variables and validate the approach in a forecasting study for the day-ahead electricity market in Germany. Our paper is one the first to tackle the issue of truly multivariate probabilistic electricity price forecasting. The contribution of this paper is therefore threefold:
  • 1.
    Methodological Contribution: We develop a regularized online estimation method for multivariate distributional regression based on the univariate work by Hirsch et al. (2024). Our algorithm allows for two layers of regularization:
    • (a)
      By leveraging online coordinate descent and LASSO-type penalties for each individual distribution parameter, we allow for high-dimensional covariate spaces.
    • (b)
      By exploiting structure in scale matrix of the multivariate distribution, we develop a path-based estimation along increasing complex dependency structures, allowing for parsimonious estimation and early stopping. We validate the trade-off between model complexity and predictive accuracy in our application study.
    Our algorithm is generic and can be applied to any parametric multivariate distribution, as long as the (log-)likelihood function and its derivatives are available. We implement the multivariate normal and $t$-distributions with three different parameterizations of the scale matrix. Further distributions can be easily added.
  • 2.
    Applied Contribution: We apply the method to probabilistic forecasting of the joint distribution of spot electricity prices. We benchmark multivariate distributional regression models to LASSO models, conformal prediction, GARCH, univariate distributional regressions approaches, partly combined with Copula constructions.
    • (a)
      We provide the first comprehensive study on multivariate probabilistic forecasting of day-ahead electricity prices, evaluating the full multivariate distribution through proper scoring rules such as the Variogram score (VS), energy score (ES), Dawid-Sebastiani score (DSS) and Log-Score (LS) for a wide range of models and including the volatile years 2021 to 2023 in our test set.
    • (b)
@@
      We show that the multivariate distributional regression, which allows modeling all distributional parameters, i.e. the mean, but also the dependence structure, conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only respectively keeping a static/unconditional dependence structure. Furthermore, we discuss the need for multivariate distributional forecasts for accurate prediction bands of the 24-hour path of day-ahead prices.
    • (c)
      We showcase the computational advantage of online estimation by benchmarking with repeated batch estimation on various re-estimation frequencies. These results provide an “efficient frontier” of computational costs against predictive accuracy.
  • 3.
@@
    Software: We provide a high-performing Python implementation using just-in-time compilation and providing a familiar, scikit-learn-like API to facilitate the usage of our package for other researchers. We contributed our code to the ondil package by Hirsch et al. (2024).
Thereby, our contribution is valuable both from a methodological and applied perspective in research and industry alike. Reproduction code can be found on GitHub. [^1] See: https://github.com/simon-hirsch/online-mv-distreg.
@@
The remainder of the paper is structured as usual: The following main Section 2 introduces the multivariate, online, regularized distributional regression model. Section 3 introduces the forecasting study, the used data and discusses multivariate forecast evaluation in detail. Section 4 presents our results and the computational costs. Section 5 concludes the paper.
@@ 2 Online Multivariate Distribution Regression
The following subsections introduce the building blocks for the multivariate, online distributional regression algorithm. The general structure of the algorithm and also the flow of the following section is outlined in Figure 3. First, we like to introduce some intuition to distributional regression. Distributional regression or “regression beyond the mean” (Kneib et al., 2023) aims at modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditioned on explanatory variables. Generally, we aim to model
$\bm{\mathrm{y}}\sim\mathcal{F}(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})\quad\text{where}\quad g_{k}(\bm{\mathrm{\theta}}_{k})=\bm{\mathrm{X}}_{k}\bm{\mathrm{\beta}}_{k}$
that is, we model the parameters $\bm{\mathrm{\theta}}_{k}$ of the distribution $\mathcal{F}$ of the response variable $\bm{\mathrm{y}}$ as regression based on the covariates in $\bm{\mathrm{X}}$. The link function $g(\cdot)$ ensures that the distribution parameters are in their domain. Naturally, this distributional model is aligned with our goal of probabilistic forecasting.
Figure 3: Online Distributional Regression. For univariate distributions (top panel), online distributional regression has been introduced by (Hirsch et al., 2024) by a combination of the iteratively reweighted least squares algorithm with online coordinate descent. We extend the approach to multivariate distributions by utilizing unconstrained scale matrix parameterizations and introducing a path-based regularization along increasingly complex dependency structures.
Section 2.1 gives a formal introduction to the univariate and multivariate case. To achieve an unconstrained estimation, we employ different parameterizations of the scale matrix in the multivariate case (Section 2.2). For the univariate case, a regularized online estimation has been introduced by Hirsch et al. (2024) based on a combination of the iteratively reweighted least squares algorithm for the estimation of distributional regression models (Section 2.3) and online coordinate descent (Section 2.4). We introduce our main contribution, the multivariate extension of this algorithm in Section 2.5 and further develop a path-based estimation along a sequence of the increasingly complex scale matrices, which allows for a regularized estimation and early stopping in Section 2.6.